# LIFT-GS: Cross-Scene Render-Supervised Distillation for 3D Language Grounding

### v1

@@ 3.2 Training 3D with 2D Supervision
Reconstruction losses: $\mathcal{L}_{\text{RGB}}$ is the 2D photometric loss, formulated as a combination of $\mathcal{L}_{1}$ loss and SSIM loss (Wang et al., 2004):
@@
Grounding losses: The grounding losses, $\mathcal{L}{\text{mask}}$, measure the prediction quality between the rendered 2D masks, $\tilde{\mathbf{M}}_{\text{2D}}$, and the ground-truth masks, $\mathbf{M}_{\text{2D}}$. Since the predicted and ground-truth masks differ in number, we employ Hungarian matching to compute the set-to-set loss. $\sigma(i)$ is the index of predicted masks matching to $i$-th ground truth mask according to the matching distance $\mathbf{d}_{\text{match}}$.
The mask shape is supervised with $\mathcal{L}_{\text{mask}}$, the combination of Focal (Lin et al., 2017) and Dice (Sudre et al., 2017) loss used in SAM (Kirillov et al., 2023):
$\mathcal{L}_{\text{CE}}$ is the correspondence loss, encouraging correct correspondences between predicted masks and language tokens. Specifically, the ground-truth mask set $\mathbf{M}_{\text{2D}}$ consists of $K$ binary masks, each corresponding to a span $\mathbf{Q}_{k}\subseteq\mathbf{Q}$ of language query embeddings $\mathbf{Q}$. $\mathbf{C}_{\sigma(i)}$ represents the correspondence probability of the $\sigma(i)$-th mask proposal across spans of the $|Q|$ language embeddings. We optimize this using a cross-entropy loss.
Feature loss: $\mathcal{L}_{\text{feat}}$ is the feature rendering loss which measures the difference between rendered feature map $\tilde{\mathbf{F}}_{\text{2D}}$ and ground-truth feature map $\mathbf{F}_{\text{2D}}$. As in (Zhu et al., 2023b), LIFT-GS uses a contrastive loss.
@@
$\tilde{\mathbf{f}}{(u,v)}$ is the feature vector of $\tilde{\mathbf{F}}_{\text{2D}}$ at location $(u,v)$. ${\mathbf{f}}_{j}$ denotes a batch of unique feature vectors from $\mathbf{F}$, and $\mathbf{f}k$ is the ground-truth feature vector corresponding to $\tilde{\mathbf{f}}{(u,v)}$.
@@ 3.3 SAM-CLIP 2D Pseudo-Label
Although LIFT-GS doesn’t need 3D annotations, getting high-quality 2D supervision remains challenging. We show how leveraging 2D foundation models for pseudo-label generation enables reasonable zero-shot performance and significantly enhances downstream tasks after fine-tuning.
As shown in Figure 3, we leverage 2D foundation models to generate pseudo-labels, including pseudo language queries and corresponding 2D masks. For each image, we apply SAM (Kirillov et al., 2023) to obtain segmentation masks. For each segmented region, we extract a CLIP (Radford et al., 2021) image embedding as the pseudo language query embedding. Since CLIP’s text and image embeddings share the same feature space, LIFT-GS can take text embeddings as input during inference. During pretraining, we concatenate these CLIP embeddings to form $\mathbf{Q}$ and construct $\mathbf{C}$, using the corresponding 2D masks as ground truth.
@@
Trained with 2D pseudo-labels, LIFT-GS can perform zero-shot 3D grounding using real text queries without fine-tuning, as shown in Figure 5. However, the zero-shot model suffers from low accuracy and struggles with complex expressions, a common limitation of CLIP-based methods that function as bag-of-words models (Yuksekgonul et al., 2023). Future improvements in pseudo-labeling, such as captioning (Meta AI, 2024) and 2D language grounding models (Liu et al., 2023), could alleviate the need for fine-tuning altogether. In this work, we focus on fine-tuning.
Overall, LIFT-GS shows that even simple pseudo-labeling strategies can be effectively distilled into 3D models. Our experiments show that pretraining with 2D pseudo-labels substantially boost performance on downstream tasks.
3.4 Architecture
LIFT-GS  is network-agnostic, imposing few architectural constraints and being readily adapted to other architectures. We describe the network used in experiments, as in Figure 4.
Encoder: Predicting 3DGS from Point Cloud
Encoder takes point clouds as input and predicts 3D Gaussian splatting (without masks) and per-point features as latent variables. Generally, the number of predicted Gaussians $|G|$ can differ from the number of input points $|P|$.
However, our focus is on demonstrating the effectiveness of differentiable rendering for structured prediction tasks like Promptable 3D segmentation. So in our experiments, we set $|G|=|P|$ with a bijective mapping to ensure consistency with point cloud-based evaluation tasks. For comparison to prior work (Zhu et al., 2023b) , LIFT-GS uses a SparseConv UNet (Contributors, 2022) as the Encoder Backbone to get per-point features, and uses an MLP as Gaussian Head to regress Gaussian parameters of $\mathbf{G}$ with features $\mathbf{F}$.
Decoder: Mask Decoder for 3D VLG
The Decoder takes per-point features from the Encoder and language embeddings $\mathbf{Q}$ as input to predict the 3D mask $\mathbf{M}\in\mathbb{R}^{N\times m}$ and the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$. The predicted $\mathbf{M}$ serves as a renderable feature of the 3DGS $\mathbf{G}$ and is subsequently used to generate 2D masks. While this work focuses on 3D vision-language grounding (3D VLG), the proposed framework is a general pipeline that can be extended to other 3D tasks with renderable outputs.
We adopt a 3D variant of the Transformer-based mask decoder in MaskFormer (Cheng et al., 2021a; Jain et al., 2025), inspired by the success of MDETR (Kamath et al., 2021) and BUTD (Jain et al., 2021). As illustrated in Figure 4, the 3D Mask Decoder Transformer takes three inputs: visual tokens from per-point features, language tokens from $\mathbf{Q}$, and $m$ learnable mask proposal tokens. It outputs $m$ predicted mask embeddings and a correspondence matrix $\mathbf{C}$. The 3D masks $\mathbf{M}$ are computed as the dot product between visual tokens and output mask embeddings. These similarity scores are then differentiably rendered into mask images $\tilde{\mathbf{M}}_{\text{2D}}$.
4 Experiments
@@
In this section, we first provide training details and demonstrate how pretraining significantly enhances downstream task performance through a set of carefully designed ablations. Additionally, we uncover several intriguing findings by scaling pretraining and fine-tuning data, as well as exploring the impact of different 2D foundation models.
@@ 4.1 Training Details
We provide details below with more details in Appendix.
Training Datasets We use ScanNet (Dai et al., 2017) as the primary dataset for downstream task fine-tuning and evaluation, as its annotations form the basis for established benchmarks. Our method’s flexibility enables training on diverse 3D datasets without requiring language annotations or instance masks. For pretraining, we primarily use ScanNet (Dai et al., 2017) and ScanNet++(Yeshwanth et al., 2023) and also explore Taskonomy(Zamir et al., 2018) and Aria Synthetic (Somasundaram et al., 2023).
Architecture Our pipeline imposes no constraints on the 3D architecture or inductive biases, allowing flexibility in model selection. For experiments, we use SparseConv UNet as the encoder backbone, a widely adopted architecture for point cloud processing. The Mask Decoder is an 8-layer Transformer decoder with a hidden state size of 512.
Training Details We train the Encoder and Decoder jointly using a batch size of 32 for 76k steps on 32 A100 GPUs. We use AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 1e-4 and weight decay of 1e-4.
@@ 4.2.2 Grounding Complex Phrases in 3D
LIFT-GS is designed to be practical. To ensure a more realistic evaluation for embodied applications, we introduce two modifications: (1) we predict 3D masks without assuming the availability of ground-truth 3D bounding boxes, and (2) we utilize sensor point clouds (Sensor PC) from RGB-D scans instead of mesh-derived point clouds (Mesh PC).
Existing methods are typically trained on mesh-derived point clouds (Mesh PC), requiring creating scene meshes from RGB-D scans, processing, and sampling point clouds. They generally rely on segments derived from clustering centers generated during human annotation. Since such information is unavailable in real-world applications, we instead use raw sensor point clouds (Sensor PC) from RGB-D scans. This setting is inherently more challenging, as reflected in the significant performance drop of BUTD-DETR (Jain et al., 2021) when transitioning from Mesh PC to Sensor PC (Table 2), consistent with findings in (Jain et al., 2024).
@@ 4.3 Ablation and Analyses
Compare to SOTA pretraining methods We compare to Ponder-v2 (Zhu et al., 2023b), a state-of-the-art method for point cloud pretraining. Ponder-v2 takes point clouds as input and predicts voxel grids, which are then used to render 2D images via NeRF (specifically NeuS, which additionally models surfaces (Wang et al., 2021)). Pretraining is supervised by 2D photometric loss and CLIP features, where features are computed based on per-pixel text labels for the mask (GT category labels and masks for each instance).
We found that the official Ponder-v2 checkpoint fails to improve performance on 3D referential grounding, likely due to the limited size and diversity of its per-pixel text labels. Therefore, we retrain the Ponder-v2 while using the SAM-CLIP pseudolabels described in Section 3.3. Training on the pseudolabels (indicated by Ponder-v2$\dagger$ in Table 3) significantly improves Acc@$0.25$ +4.5% vs. the official Ponder-v2, and +3.2% vs. training from scratch. Since both approaches use render-supervision, this experiment underscores the advantages of pseudolabel distillation in data-scarce regimes. Moreover our method does not rely on human-annotated per-pixel text labels, making it highly scalable to other datasets, as shown in Table 6.
@@
Loss Ablation Existing pretraining pipelines primarily focus on the encoder (Zhu et al., 2023b; Banani et al., 2021), whereas the render-supervised formulation can pretrain the entire architecture in a unified manner. To assess the impact of the grounding loss, which applies only to the decoder, we compare models pretrained with different losses in Table 4.
A model trained with $\mathcal{L}_{\text{ground}}$ alone (row 2) still substantially improves downstream task performance, performing only slightly worse than the model trained with the full loss (row 5), demonstrating the effectiveness of the grounding loss. Furthermore, comparing models with and without $\mathcal{L}{\text{ground}}$ (row 5 vs. row 4) clearly shows that $\mathcal{L}_{\text{ground}}$ significantly enhances downstream performance, particularly in more challenging scenarios (IoU thresholds of 0.5 and 0.75).
@@ 4.4 Data Scaling
Finetuning Data Scaling We limit the fine-tuning data size to $0.1$, $0.2$, and $0.5$ of the full dataset for 3D referential grounding to analyze performance variation with different fine-tuning ratios. The results are presented in Figure 6 and Table 5, with additional findings provided in Appendix A.3.
Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data. For instance, the pretrained model fine-tuned on $20\%$ of the data achieves performance comparable to training from scratch with $50\%$ of the data. The benefits of pretraining are even more pronounced at higher IoU thresholds. Overall, a pretrained model can achieve the same performance as a model trained from scratch on all the data using only $30{-}40\%$ of the fine-tuning data.
This observation matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021). And interestingly, the fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, which underscores that 3D VLG models are currently operating in the data scarce regime.
@@
Pretraining Data Scaling A key advantage of our framework is its ability to pretrain on large-scale RGB-D scans without requiring 3D or language annotations, ensuring scalability. We scale the pretraining dataset and analyze its impact on fine-tuning performance, as shown in Table 6.
[…]
Since our method doesn’t need 3D masks or language annotations, expanding the pretraining dataset with real-world scans is fairly easy. We anticipate continuous improvements as larger-scale data collections become available.
@@
Weaker CLIP and SAM The main experiments use SAM-H and CLIP-L for pseudo-labeling. Replacing them with smaller models, MobileSAM (Zhang et al., 2023a)(ViT-tiny)  and CLIP-B, leads to a noticeable performance drop, especially at higher accuracy thresholds. This suggests that our model benefits from advancements in 2D foundation models, highlighting the importance of stronger 2D models.
@@
Captions from VLMs We explore an alternative pseudo-labeling strategy by combining SAM with the LLAMA-V model (details in Appendix). After segmenting objects in 2D images using SAM, we prompt LLAMA-V to describe the segmented regions and encode these descriptions into CLIP embeddings.
Pretraining with this pseudo-labeling method achieves performance comparable to our original pipeline with SAM-H and CLIP-L, highlighting the flexibility and extensibility of our framework. We believe text-based captions hold significant potential for future research, and our pipeline is well-positioned to benefit from advancements in this area.
@@ 5 Conclusion
We present LIFT-GS, a model-agnostic pipeline for training 3D VLG models without 3D supervision by leveraging Gaussian Splatting and 2D foundation models. Beyond enabling zero-shot 3D grounding, LIFT-GS significantly improves downstream performance when fine-tuned with limited 3D data. It also scales effectively, benefiting from larger datasets and stronger 2D foundation models, suggesting continuous improvements with increased data. Limitations and future directions are discussed in the Appendix.

### v2

@@ 3.2.2 Per-Pixel Losses
Reconstruction loss, $\mathcal{L}_{\text{RGB}}$: supervises photometric reconstruction using $L_{1}$ and SSIM losses (Hong et al., 2024; Zhu et al., 2023b; Wang et al., 2004)
@@
Feature loss, $\mathcal{L}_{\text{feat}}$: uses CLIP-style contrastive regularization to align rendered features $\tilde{\mathbf{F}}_{\text{2D}}$ with ground-truth features $\mathbf{F}_{\text{2D}}$ as in (Zhu et al., 2023b):
@@
where $\tilde{\mathbf{f}}_{(u,v)}$ is the rendered feature at pixel $(u,v)$ from $\tilde{\mathbf{F}}_{\text{2D}}$, $\mathbf{f}_{k}$ is the corresponding ground-truth feature, and $\mathbf{f}_{j}$ represents the batch of unique features.
3.3 Architecture
LIFT-GS is network-agnostic with minimal architectural constraints. The design can be applied to other architectures, with constraints only arising from the specific losses used. This section describes the network used in our experiments, shown in Figure 5.
3.3.1 Grounding Decoder: (see Grounding Losses)
The grounding decoder is a transformer based on MaskFormer (Cheng et al., 2021b) that predicts the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$ and 3D Gaussian masks $\mathbf{M}\in\mathbb{R}^{N\times m}$. LIFT-GS uses the mask decoder from UniVLG (Jain et al., 2025) with minimal modifications.
The transformer takes Gaussian features $\mathbf{G}$ and language query embeddings $\mathbf{Q}$ as inputs, using a set of learnable tokens as 3D mask proposals. Cross-attention is computed between the learnable tokens and both the language and Gaussian tokens. After extensive information exchange within the transformer, $\mathbf{M}$ (or $\mathbf{C}$) is computed as dot products between mask proposal tokens and Gaussian tokens (or mask proposal tokens and language tokens, respectively). The detailed LIFT-GS architecture and hyperparameter settings used in our experiments are provided in Appendix A.
3.3.2 Gaussian Decoder Head: (see Per-Pixel Losses)
LIFT-GS predicts $\mathbf{G}\in\mathbb{R}^{N\times F}$ using a learned pointwise MLP applied to the point cloud encoder outputs. While the number of predicted Gaussians $|G|$ can differ from input points $|P|$, we set $|G|=|P|$ with a bijective mapping for consistency with point cloud evaluation tasks. Notably, the Gaussian decoder does not require direct supervision–3D Gaussians can be treated as latent variables for 3D VLG, as shown in Table 4 when per-pixel losses are disabled.
3.3.3 Input Encoders:
Pointcloud Encoder: Tokenizes RGB point clouds of shape $|P|{\times}6$ (xyz + RGB) using a sparse convolutional UNet (Contributors, 2022), following PonderV2 (Zhu et al., 2023b). Weights are randomly initialized and learned.
Text Encoder: LIFT-GS uses CLIP text embeddings, which remain frozen during training.
These encoders represent common choices in 3D VLG. Stronger architectural choices like transformer-based pointcloud encoders or different text embeddings would likely improve performance.
@@ 3.4 SAM-CLIP 2D Pseudo-Label
While LIFT-GS eliminates the need for 3D annotations, obtaining high-quality 2D supervision remains challenging. We show one way in which 2D foundation models can generate pseudo-labels that enable reasonable zero-shot performance and significantly enhance downstream fine-tuning.
As shown in Figure 4, we generate pseudo-labels using SAM (Kirillov et al., 2023) and CLIP (Radford et al., 2021). For each image, SAM provides segmentation masks, and for each segmented region, we extract CLIP image embeddings as pseudo language query embeddings. Since CLIP’s text and image embeddings share the same feature space, LIFT-GS can use text embeddings during inference. We concatenate these CLIP embeddings to form $\mathbf{Q}$ and construct $\mathbf{C}$ s.t. each instance maps to exactly one query token.
@@
With 2D pseudo-labels, LIFT-GS performs zero-shot 3D grounding using real text queries without fine-tuning (Figure 6). However, zero-shot performance suffers from low accuracy and struggles with complex expressions – a common limitation of CLIP-based methods that function as bag-of-words models (Yuksekgonul et al., 2023). Future improvements in pseudo-labeling, such as better captioning (Meta AI, 2024) and 2D language grounding models (Liu et al., 2023), could reduce reliance on fine-tuning.
LIFT-GS demonstrates that even simple pseudo-labeling strategies can be effectively distilled into 3D models. In the experiments below, pretraining with 2D pseudo-labels substantially improves downstream task performance. Fine-tuning data scaling results are consistent with established transfer learning “scaling laws” (Hernandez et al., 2021).
@@
4 Experiments
Although supervised via 2D loss, LIFT-GS is a fully 3D model that explicitly outputs 3D masks. Beyond the zero-shot setting using 2D pseudo-labels (Figure 6), LIFT-GS can be readily fine-tuned with 3D annotation data using 3D losses for significantly improved performance. To enhance practical applicability, we focus on fine-tuning with 3D annotations and demonstrate how 2D model distillation boosts performance.
In this section, we first provide training details and show how pretraining significantly improves downstream task performance through a series of carefully designed ablations. We also reveal several insights from scaling the amount of pretraining and fine-tuning data, and explore the impact of different 2D foundation models.
@@ 4.1 Training Details
We provide details below with more details in the Appendix.
Datasets We use ScanNet (Dai et al., 2017) as the primary dataset for downstream task fine-tuning and evaluation, as its annotations form the basis for established benchmarks. We primarily pretrain using ScanNet (Dai et al., 2017) for comparison to other methods. LIFT-GS enables training on diverse unlabeled 3D datasets, and we show additional pretraining scaling experiments using ScanNet++(Yeshwanth et al., 2023), Taskonomy (Zamir et al., 2018) and Aria Synthetic (Somasundaram et al., 2023).
Architecture LIFT-GS method imposes minimal architectural constraints. In the following experiments, backbones consist of a text encoder (frozen CLIP-L), point cloud encoder (Sparse 3D UNet (Çiçek et al., 2016)), and grounding decoder (8-layer Transformer, hidden size 512 following (Jain et al., 2025)), and an MLP for the Gaussian decoder.
Training Models are trained end-to-end for 76k steps with batch size 32 on 32 A100s using AdamW (Loshchilov & Hutter, 2017) (lr=1e-4, weight decay=1e-4). Point clouds are voxel-downsampled to 5cm for an average of  50k points, and we render masks at resolution $512{\times}512$ to ensure small masks are captured. Complete implementation details are provided in the appendix.
@@ 4.2.2 Grounding Complex Phrases in 3D
LIFT-GS is designed to be practical and we evaluate it using the “real-world” settings used in more recent 3D VLG work where (1) we predict 3D masks without assuming known ground-truth 3D bounding boxes, and (2) we utilize sensor point clouds (Sensor PC) from RGB-D scans instead of using mesh-derived point clouds that leak label information (Mesh PC). This realistic setting is more challenging, as reflected in the significant performance drop of BUTD-DETR (Jain et al., 2021) when transitioning from Mesh PC to Sensor PC (Table 2), consistent with findings in (Jain et al., 2024). A more complete comparison of these settings is provided in (Jain et al., 2021, 2025; Arnaud et al., 2025).
@@ 4.3 Pretraining Ablations
Compare to SOTA pretraining methods We compare against PonderV2 (Zhu et al., 2023b), a state-of-the-art point cloud pretraining method that also uses render-supervision. Since the official PonderV2 relies on limited human-annotated text labels, we retrain it using our SAM-CLIP pseudo-labels for fair comparison (PonderV2$\dagger$ in Table 3). This demonstrates the value of distillation, improving the performance over using GT labels (45.4% vs 40.9% Acc@0.25). Moreover, LIFT-GS substantially outperforms PonderV2$\dagger$ (47.5% vs 45.4% Acc@0.25), underscoring the impact of multimodal decoder architectures enabled by the LIFT-GS render-supervised formulation.
@@
Loss Ablation Existing pretraining pipelines primarily focus on the encoder (Zhu et al., 2023b; Banani et al., 2021), whereas the render-supervised formulation can pretrain the entire architecture in a unified manner using the grounding loss. We find that grounding loss alone can be used to pretrain the model end-to-end in Table 4. A model trained with $\mathcal{L}_{\text{ground}}$ alone (row 2) substantially improves downstream task performance, performing only slightly worse than the model trained with all losses (row 5). Furthermore, comparing models with and without $\mathcal{L}{\text{ground}}$ (row 5 vs. row 4) clearly shows that $\mathcal{L}_{\text{ground}}$ significantly enhances downstream performance, particularly in more challenging scenarios (IoU thresholds of 0.5 and 0.75).
@@ 4.4 Data Scaling
LIFT-GS exhibits strong scaling properties that reveal 3D VLG operates in a severely data-scarce regime.
Finetuning Data Scaling We observe that pretraining effectively ”multiplies” the fine-tuning dataset by approximately 2x. As shown in Figure 7 and Table 5, a pretrained model using $50\%$ of fine-tuning data matches the performance of training from scratch with $100\%$ data. This scaling coefficient remains constant across different data amounts ($10\%$, $20\%$, and $50\%$) without diminishing–matching empirical scaling laws from other modalities (Hernandez et al., 2021). The benefits are most pronounced at higher IoU thresholds, where a pretrained model achieves scratch-level performance using only $30{-}40\%$ of the fine-tuning data. Additional results on Instance Segmentation are provided in Appendix A.3.
@@
Pretraining Data Scaling Expanding pretraining data consistently improves downstream performance (Table 6). Adding ScanNet++ yields notable gains (+0.8%), while incorporating Taskonomy and Aria Synthetic provides additional improvements despite distribution differences (likely due to the mesh reconstruction quality in Taskonomy).
Comparing the scaling results from pretraining and fine-tuning demonstrates the strong data efficiency of LIFT-GS. As shown in Table 6, adding ScanNet++ ( 30% of the data used for pretraining) yields a performance gain equivalent to adding 15% more ScanNet fine-tuning data with 3D annotations, based on the curve in Figure 7. This indicates an effective transfer ratio of roughly 1:2—i.e., collecting twice as many raw videos provided improvements comparable to building a fully annotated 3D SFT dataset.
Therefore, pretraining on ScanNet++ is not only highly effective but also cost-efficient, especially considering that annotating 3D referential grounding data requires significantly more effort than collecting raw videos alone.
Data Scarcity Implications The consistent 2× multiplier without saturation, combined with continued gains from more pretraining data, strongly suggests that current 3D VLG models are severely limited by data availability–opening a path for future 3D VLG improvements through scaling alternate sources (such as 2D foundation models).
@@
Weaker CLIP and SAM The main experiments use SAM-H and CLIP-L for pseudo-labeling. Replacing them with smaller models, MobileSAM (Zhang et al., 2023a)(ViT-tiny)  and CLIP-B, leads to a noticeable performance drop, especially at higher accuracy thresholds. This suggests that render-supervised distillation directly benefit from advancements in 2D foundation models.
@@
Captions from LMMs Table 7 shows results using large LMMs instead of CLIP to generate queries (LLAMA-3V + SAM grounding, details in Appendix). After segmenting objects in 2D images using SAM, we prompt LLAMA-V to describe the segmented regions. Pretraining with these captions achieves performance comparable to our original pipeline with SAM-H and CLIP-L. As LMMs continue to improve, we believe text-based captions hold significant potential for future research, and the approach highlighted in the paper is positioned to benefit from LMM improvements.
@@ 5 Conclusion
LIFT-GS tackles data scarcity that limits 3D VLG by introducing render-supervised distillation from 2D VLM models. By training 3D models using only 2D supervision from models like SAM and CLIP, LIFT-GS achieves state-of-the-art performance for 3D VLG. Our findings, including consistent 2x data multiplication effects, reveal that 3D grounding currently operates with substantial data limitations. LIFT-GS circumvents a key 3D annotation bottleneck by introducing a scalable training approach that benefits directly from advancements in frontier multimodal language models. It offers a practical technique to leverage progress in 2D to accelerate the development of other data-scarce capabilities essential for robotics, AR/VR, and embodied AI.
Acknowledgment
This work is done during Ang Cao’s internship at Meta. We thank Andrew Owens, Andrea Vedaldi, Stella Yu, Ziyang Chen, Yiming Dou, Xuanchen Lu for their helpful discussion and feedback.
Impact Statement
We propose a method for training 3D models without 3D supervision, advancing 3D vision-language research. Our approach significantly improves 3D referential grounding, a key task for robotics, embodied AI, and AR/VR applications. The resulting model enables agents to precisely localize objects from language inputs, bridging high-level reasoning with real-world actions.
As our model is distilled from 2D foundation models, it may inherit their biases. However, since our primary task is grounding, it is unlikely to introduce significant aesthetic biases.